OpenStudy (anonymous):

Let f(x) = -4x + 7 and g(x) = 10x - 6, find f(g(x)). -40x + 24 -40x + 31 -40x + 64 -40x + 70

OpenStudy (anonymous):

its b

OpenStudy (anonymous):

always start in the middle \[f(g(x))=f(10x-6)\]i.e. replace \(g(x)\) by \(10x-6\)

OpenStudy (anonymous):

then where you see an \(x\) in \(f(x)\) replace that by \(10x-6\) i.e. \[f(\spadesuit)=-4\spadesuit +7\] so \[f(10x-6)=-4\times (10x-6)+7\] then some algebra to clean it up

OpenStudy (anonymous):

you get \[-4(10x-6)+7=-40x+24+7=-40x+31\]

OpenStudy (anonymous):

Thank you so much for explaining! I get it know! :)

OpenStudy (anonymous):

yw